Given that U is a function of x, y and z and A a vector field, prove that:

∇⋅(UA)=(∇U)⋅A+U(∇⋅A).



1
Expert's answer
2021-12-14T09:55:15-0500

"∇⋅(UA)=\frac{\partial}{\partial x}(UA_x)+\frac{\partial}{\partial y}(UA_y)+\frac{\partial}{\partial z}(UA_z)="


"=\frac{\partial U}{\partial x}(A_x)+\frac{\partial U}{\partial y}(A_y)+\frac{\partial U}{\partial z}(A_z)+\frac{\partial A_x}{\partial x}(U)+\frac{\partial A_y}{\partial y}(U)+\frac{\partial A_z}{\partial z}(U)="


"=(\u2207U)\u22c5A+U(\u2207\u22c5A)"